Routine check for DGS maintenance day (12/03)
Check if the power supply can be controlled by BIO.
Check if the script can be started.
Confirmed that the limit switch installed is in Green.
However, ETMX is in use, so not confirmed yet.
Confirmed that the Release Sensor is 0. 1: Seated
ETMX, ETMY, ITMX, ITMY : No problem
ETMX_F1_1 : Abnormality disappeared as of last week (Nov.25 10:32:50 JST).
They are replacing the flip cable.
No problem (all NGs are known)
=> Confirmation only, no recovery work was done.
PCAL(EX1,EX2,EY1,EY2):Not confirmed since it is ON only when used.
Check if the temperature acquisition is working properly.